40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 286 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  2965   Fri May 21 08:28:29 2010 steveConfigurationVACTP3 turned off

Quote:

[Jenne, Kiwamu, and Steve via phone]

Around 9:30pm, Kiwamu and I came back from dinner, and were getting ready to begin the beam scan measurements.  I noticed that one of the vacuum pumps was being very loud.  Kiwamu noted that it is the fore pump for TP3's turbo, which he and Steve replaced in January (elog 2538).  We had not noticed these noises before leaving for dinner, around 8pm.

We called Steve at home, and he could hear the noise through the phone. He said that even though it was really loud, since it was reading 3.3mTorr (on the display of the controller, in the vacuum rack just above head-height) which is close to the nominal value, it should be fine to leave.  He will check it out in the morning.  If it had been reading at or above ~1Torr, that's indicative of it being really bad, and we would have needed to shut it off.

For future reference, in case we need to turn it off, Steve said to use the following procedure:

1. Close VM3, to isolate the RGA, which is what this pump is currently (while we're at atmosphere) pumping on.  I don't know if there are other things which would need to be shut at this stage, if we were at vacuum nominal.

2. Close VM5, which is right in front of TP3, so TP3's pump is just pumping on itself.

3. Push the "Stop" button on the Turbo controller for TP3, in the vacuum rack, about waist level.  Turning off the turbo will also turn off the fore pump.

UPDATE, 1am: The controller in the rack is reading 3.1mTorr, so the pump, while still noisy, still seems to be working.

 The foreline pressure of TP3 is 2.9 mTorr  The drypump is loosing it's bearing and it is very noisy.

V3, V5 closed and TP3 small turbo controller off. This turned off the noisy forepump that has to be replaced.

RGA is running at cc4 2e-6 Torr

The RGA was turned off at cc4 1e-5 Torr

  2964   Fri May 21 00:51:06 2010 JenneUpdateIOOFirst MC mode measuring (hopefully) done

[Jenne, Kiwamu, Steve]

Round 1 of measuring the MC mode is pretty much done.  Yay.

Earlier today, Steve and I launched the MC beam off the flat mirror just after the Faraday, and sent it down toward ETMY(new convention). We ended up not being able to see it all the way at the ETM because we were hitting the beam tube, but at the ITM chamber we could see that the beam looked nice and circle-y, so wasn't being clipped in the Faraday or anywhere else.  To do this we removed 2 1inch oplev optics.  One was removed from the BS table, and wrapped in foil and put in a plastic box.  The other was just layed on its' side on the BS table. 

I then took the beam out of the BS chamber, in order to begin measuring the mode.  I left the flat fixed mirror in the place of what will be PZT SM1, and instead used the PZT mirror to turn the beam and get it out the BS chamber door.  (Thoughts of getting the beam to the BS oplev table were abandoned since this was way easier, since Kiwamu and Steve had made the nifty table leg things.)  Kiwamu and I borrowed an 2inch 45P Y1 optic from the collection on Koji's desk (since we have ZERO 2inch optics on the random-optic-shelf....no good), to shoot the beam down the hallway of the Yarm (new convention).  We used the beam scan on a rolling cart to measure the beam at various distances.  I made some sweet impromptu plum bobs to help make our distance measurements a bit more accurate.

We stopped at ~25 feet from the BS chamber, since the spot was getting too big for the beam scanner.  If it turns out that I can't get a good fit with the points I have, I'll keep everything in-chamber the same, and do the farther distances using the good ol' razor blade technique.

I have measurements for the distances between the beam scan head and the opening of the BS chamber.  Tomorrow, or very soon after, I need to measure the distances in-chamber between the MC and the BS chamber opening.  Plots etc will come after I have those distances.

Next on the to-do list:

1.  Measure distances in-chamber for first mode scan.

2.  Plot spot size vs. distance, see if we need more points. Take more points if needed.

3.  Put in MMT1, repeat measurements.

4.  Put in MMT2, rinse and repeat.

5.  Move the PZT mirror to its new place as SM1, and figure out how to connect it.  Right now the little wires are hooked up on the BS table, but we're going to need to make / find a connector to the outside world from the IOO table. This is potentially a pretty big pain, if we don't by happenstance have open connectors on the IOO table.

  2963   Fri May 21 00:30:44 2010 JenneUpdateVACTP3 fore pump is very loud

[Jenne, Kiwamu, and Steve via phone]

Around 9:30pm, Kiwamu and I came back from dinner, and were getting ready to begin the beam scan measurements.  I noticed that one of the vacuum pumps was being very loud.  Kiwamu noted that it is the fore pump for TP3's turbo, which he and Steve replaced in January (elog 2538).  We had not noticed these noises before leaving for dinner, around 8pm.

We called Steve at home, and he could hear the noise through the phone. He said that even though it was really loud, since it was reading 3.3mTorr (on the display of the controller, in the vacuum rack just above head-height) which is close to the nominal value, it should be fine to leave.  He will check it out in the morning.  If it had been reading at or above ~1Torr, that's indicative of it being really bad, and we would have needed to shut it off.

For future reference, in case we need to turn it off, Steve said to use the following procedure:

1. Close VM3, to isolate the RGA, which is what this pump is currently (while we're at atmosphere) pumping on.  I don't know if there are other things which would need to be shut at this stage, if we were at vacuum nominal.

2. Close VM5, which is right in front of TP3, so TP3's pump is just pumping on itself.

3. Push the "Stop" button on the Turbo controller for TP3, in the vacuum rack, about waist level.  Turning off the turbo will also turn off the fore pump.

UPDATE, 1am: The controller in the rack is reading 3.1mTorr, so the pump, while still noisy, still seems to be working.

  2962   Fri May 21 00:21:24 2010 JenneUpdateSAFETYDon't walk down the Y(new convention) arm tonight!

All clear.

Quote:

Please don't go down the Yarm (Old Xarm) for right now, or if you do, please be very careful.  Kiwamu and I are set up to take beam scan measurements down the walkway, and so there are some cables / carts / other stuff down there.  We are going to get dinner really quickly before beginning the measurements. 

Right now, the PSL shutter is Closed, so there is no beam hazard outside of the chambers, just crowded space hazard.

 

  2961   Thu May 20 20:03:37 2010 JenneUpdateSAFETYDon't walk down the Y(new convention) arm tonight!

Please don't go down the Yarm (Old Xarm) for right now, or if you do, please be very careful.  Kiwamu and I are set up to take beam scan measurements down the walkway, and so there are some cables / carts / other stuff down there.  We are going to get dinner really quickly before beginning the measurements. 

Right now, the PSL shutter is Closed, so there is no beam hazard outside of the chambers, just crowded space hazard.

  2960   Thu May 20 14:18:59 2010 kiwamuUpdateGreen Lockingmode profile of 40m cavity

The mode profile of the green beam going through 40m cavity was measured.

According to the fitting the coupling efficiency to the cavity is 98.46%, but still the beam looks loosely focused.

This measurement has been done by using the oplev legs (entry #2957) to allow the beam to go through the 40m walkway.

With a beam scan set on a movable cabinet, I measured it along the 40m chamber.

Since the plot looks not so nice,  I am going to work on this measurement a little bit more after I improve the mode matching.

 


Here is the parameters from the fitting

target waist [mm] 2.662
measured waist x [mm] 2.839
measured waist x [mm] 3.111
   
target waist position [m] 43.498
measured waist position x [m] 42.579
measured waist position y [m] 38.351

 

I believe the error for the travel length was within 0.5 meter. The length was always measured by a tape measure.

A thing I found was that: spatial jittering of the beam gets bigger as the beam goes further. This is the main source of the error bar for the spot size. 

 

 

Attachment 1: MMT40mcavity.png
MMT40mcavity.png
  2959   Thu May 20 13:29:40 2010 kiwamuUpdateGreen Lockingmode profile at PPKTP crystal

 I measured again the mode profile of the beam going through the PPKTP crystal by using the beam scan.

The aimed beam waist is 50 um (as described in entry 2735),

and the measured profile had pretty good waist of wx=51.36 +/- 0.0999 um and wy=49.5 +/- 0.146 um 

The next things I have to do are - (1). re-optimization of the temperature of the crystal (2). measurement of the conversion efficiency

The attached figure is the result of the measurement. 

Attachment 1: PPKTPmode.png
PPKTPmode.png
  2958   Thu May 20 13:12:28 2010 josephbUpdateCDSPreparations for testing lsc,lsp, scy,spy together

In /cvs/cds/caltech/target/fb modified:

master: cleaned up so only io1 (IO processor), LSC, LSP, SCY, SPY were listed, along with their associated tpchan files.

daqdrc: fixed "dcu_rate 9 = 32768" to "dcu_rate 9 = 65536" (since the IO processor is running at 64k)

Added "dcu_rate 21 = 16384" and "dcu_rate 22 = 16384"

Changed "set gds_server = "megatron" "megatron" "megatron" 9 "megatron" 10 "megatron" 11;" to

set gds_server = "megatron" "megatron" 9 9;

The above change was made after reading Rolf's Admin guide: http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/CDS?action=AttachFile&do=get&target=RCG_admin_guide.pdf
The set gds_server is simply telling which computer the gds daemons are running on, and we don't need to do it 5 times.


In /cvs/cds/caltech/gds/params modified:

testpoint.par: added C-node7 and C-node8 for SCY and SPY respectively.

  2957   Thu May 20 12:34:46 2010 kiwamuConfiguration40m Upgradingoptical breadboards with legs

Yesterday Steve and I revived two legs to mount some optical breadboards outside of the end table.

These legs had been used as oplev's mounts many years ago, but now they are served for 40m upgrading. These are really nice.

By putting them on the side of the end table, a mirror mounted on the top of the leg can reflect the beam outside of the end table.

Once we pick off the green beam from the end table to its outside, the green beam can propagate through the 40m walkway along the Y-arm.

So that we can measure the beam profile as it propagates.

These legs are also going to be used during mode matching of the vacuum optics.

Attachment 1: leg1_small.png
leg1_small.png
Attachment 2: leg2_small.png
leg2_small.png
  2956   Thu May 20 12:10:44 2010 kiwamuUpdatePhotosETMY end table

 I updated the photo of ETMY end table on the wiki.

http://lhocds.ligo-wa.caltech.edu:8000/40m/Optical_Tables

Attachment 1: ETMY_s.png
ETMY_s.png
  2955   Thu May 20 10:06:56 2010 AidanHowToPhase CameraPhase Camera- raw data video


 

  2954   Wed May 19 22:28:05 2010 KojiUpdateIOOHooray! We locked the MC! (and some other stuff)

Good! What was the key?

The MC2 spot looks very high, but don't believe the TV image. Believe the result of script/A2L/A2L_MC2. What you are looking at is the comparison of the spot at the front surface and the OSEMs behind the mirror.

Quote:

[Jenne, Kevin]

We opened up the MC chambers again, and successfully got the MC locked today!  Hooray!  This meant that we could start doing other stuff....

First, we clamped the Faraday.  I used the dog clamps that Zach left wrapped in foil on the clean cart.  I checked with a card, and we were still getting the 00 mode through, and I couldn't see any clipping.  2 thumbs up to that.

Then we removed the weight that was on the OMC table, in the way of where MMT2 needs to go.  We checked the alignment of the MC, and it still locks on TEM00, but the spot looks pretty high on MC2 (looking at the TV view). We're going to have to relevel the table when we've got the MMT2 optic in the correct place.

We were going to start moving the PZT steering mirror from the BS table to the IOO table, place MMT2 on the OMC table, and put in a flat mirror on the BS table to get the beam out to the BS oplev table, but Steve kicked us out of the chambers because the particle count got crazy high.  It was ~25,000 which is way too high to be working in the chambers (according to Steve).  So we closed up for the day, and we'll carry on tomorrow.  

Photos of the weight before we removed it from the OMC table, and a few pictures of the PZT connectors are on Picasa

 

  2953   Wed May 19 16:09:11 2010 josephbUpdateCDSRacks to small for IO Chassis rails

So I discovered the hard way that the racks are not standard width, when I was unable to place a new IO chassis into the racks with rails attached.  The IO chassis is narrow enough to fit through without the rails however. 

I've talked to Steve and we decided on having some shelves made.  I've asked Steve to get us 6.  1 for each end (2), 1 for SUS, 1 for LSC, 1 for IO, and 1 extra.

  2952   Wed May 19 16:00:18 2010 JenneUpdateIOOHooray! We locked the MC! (and some other stuff)

[Jenne, Kevin]

We opened up the MC chambers again, and successfully got the MC locked today!  Hooray!  This meant that we could start doing other stuff....

First, we clamped the Faraday.  I used the dog clamps that Zach left wrapped in foil on the clean cart.  I checked with a card, and we were still getting the 00 mode through, and I couldn't see any clipping.  2 thumbs up to that.

Then we removed the weight that was on the OMC table, in the way of where MMT2 needs to go.  We checked the alignment of the MC, and it still locks on TEM00, but the spot looks pretty high on MC2 (looking at the TV view). We're going to have to relevel the table when we've got the MMT2 optic in the correct place.

We were going to start moving the PZT steering mirror from the BS table to the IOO table, place MMT2 on the OMC table, and put in a flat mirror on the BS table to get the beam out to the BS oplev table, but Steve kicked us out of the chambers because the particle count got crazy high.  It was ~25,000 which is way too high to be working in the chambers (according to Steve).  So we closed up for the day, and we'll carry on tomorrow. 

 

Photos of the weight before we removed it from the OMC table, and a few pictures of the PZT connectors are on Picasa

  2951   Wed May 19 14:36:46 2010 AidanHowToPhase CameraPhase Camera algorithm and stuff

 I had a think about the algorithm we might use for the phase camera measurement. MATLAB has an fft function that will allow us to extract the data that we need with a single command.

We record a series of images from a camera and put them into a 3D array or movie, image_arr, where the array parameters are [x-position, y-position, time], i.e. a 2D slice is a single frame from the camera. Then we can do an FFT on that object with the syntax, f3D = fft(image_arr, [ ], 3), which only does the FFT on the temporal components. The resulting object is a 3D array where each 2D slice is an 2D array of amplitude and phase information across the image for a single temporal frequency of the movie.

So if we recorded a movie for 1s where the sample rate is 58Hz, then the 1st frame of f3D is just a DC image of the movie, the 2nd frame are the complex 1Hz components of the movie, etc all the way up to 29Hz. 

Suppose then that we have a image, part of which is being modulated, e.g. a chopper wheel rotating at 20 or 24Hz, or a laser beam profile which contains a 1kHz beat between a sideband and a reference beam. All we have to do is sample at at least twice that modulation frequency, run the command in MATLAB, and then we immediately get an image which contains the phase and magnitude information that we're interested in (in the appropriate 2D slice o the FFT).

As an example, I recorded 58 frames of data from a camera, sampling at 58Hz, which was looking at a spinning chopper wheel. There was a white sheet of paper behind the wheel which was illuminated from behind by a flashlight. The outer ring was chopping at 24Hz and the inner ring was chopping at 20Hz. I stuck all the images into the 3D array in MATLAB, did the transformation and picked out the DC, 20Hz and 24Hz signals. The results are shown in the attached PDFs which are:

  1. phase_camera_DC_comp.pdf - a single image from the camera and the DC component (zoomed in) of the FFT
  2. phase_camera_F1_comp.pdf - the magnitude and phase information of the 20Hz component of the FFT
  3. phase_camera_F2_comp.pdf - the magnitude and phase information of the 24Hz component of the FFT (this PDF contains a typo that says 25Hz).
  4. load_raw_data.m - the MATLAB routine that loads the saved data from the camera and does the FFT

You can, and I have, run the MATLAB engine from C directly. This will allow you to transfer the data from the camera to MATLAB directly in memory, rather than via the disk, but it does need proper memory allocation to avoid segmentation faults - that was too frustrating for me in the short term. In this case, the 58 frames were recorded to a file as a contiguous block of data which I then loaded into MATLAB, so it was slower than it might've otherwise been. Also the computer I was running this on was a bit of a clunker so it took a bit of time to do the FFT.

The data rate from the camera was 58fps x (1024 x 1024) pixels per frame x 2 bytes per pixel = 116MB per second. If we were to use this technique in a LIGO phase camera, where we want to measure a modulation which is around 1kHz, then we'd need a sample rate of at least 2kHz, so we're looking at at least a 30x reduction in the resolution. This is okay though - the original phase camera had only ~4000 spatial samples. So we could use, for instance, the Dalsa Falcon VGA300 HG which can give 2000 frames per second when the region of interest is limited to 64 pixels high.

Attachment 1: phase_camera_DC_comp.pdf
phase_camera_DC_comp.pdf
Attachment 2: phase_camera_F1_comp.pdf
phase_camera_F1_comp.pdf
Attachment 3: phase_camera_F2_comp.pdf
phase_camera_F2_comp.pdf
Attachment 4: load_raw_data.m
% load a raw data file into MATLAB

fid = fopen('phase_camera_data.dat');
n1 = 750;
A3D = ones(n1, n1, 58);

for jj = 1:58
    A = fread(fid, [1024, 1024], 'uint16');
    A3D(:,:,jj) = A((512-floor(n1/2)):(512-floor(n1/2))+n1-1, ...
                    (512-floor(n1/2)):(512-floor(n1/2))+n1-1);
... 64 more lines ...
  2950   Tue May 18 23:03:08 2010 JenneUpdateIOONo real progress....

[Jenne, Kevin]

No real progress today.  We opened the chambers and again tried to lock the MC.  Gave up after ~2.5 hours (and closed up the chambers with light doors, replaced manual beam block, etc...).  With Koji's helpful coaching, hopefully we'll finally get it done tomorrow.  Then we can move forward with the actual to-do list. 

 

  2949   Tue May 18 16:44:35 2010 KojiUpdateIOOFirst steps toward MC mode measuring

Here is the upadted list http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/Optics

Quote:

I will update how the mirrors should be migrated from the table to the table. 

 

  2948   Tue May 18 16:19:19 2010 josephbUpdateCDSWe have two new IO chassis

We have 2 new IO chassis with mounting rails and necessary boards for communicating to the computers.  Still need boards to talk to the ADCs, DACs, etc, but its a start.  These two IO chassis are currently in the lab, but not in their racks.

They will installed into 1X4 and 1Y5 tomorrow.  In addition to the boards, we need some cables, and the computers need the approriate real time operating systems setup.  I'm hoping to get Alex over sometime this week to help work on that.

  2947   Tue May 18 15:09:02 2010 steveConfigurationSAFETYPSL output shutter in order again

Quote:

As we learned yesterday, the PSL laser power out put mechanical shutter is not working in the remote mode. It only works in local manual mode.

Do not rely on the MEDM screen monitor readout! The position is only changing on the monitor. The main beam must be blocked before the output periscope.

Ben found the Sorenson 5V ps off. It was turned off since our last scheduled power outage. I wonder what else is running on 5V in the PSL? This power supply should be on the

"alarm handler"  list to avoid future repeat of this condition. However a real safety switch would have confirming position sensors of the shutter open or closed. Is there such thing at the sites?

  2946   Tue May 18 14:30:31 2010 josephbUpdateCDSLSC.mdl problem found and fixed

After having checked old possibilities and deciding I wasn't imagining the lsc.mdl file not working, but working as another name, I tracked Alex down and asked for help.

After scratching our heads, we finally tracked it down to the RCG code itself, as opposed to any existing code.

Apparently, the skeleton.st file (located in /home/controls/cds/advLigoRTS/src/epics/util/) has special additional behavior for models with the following names: lsc, asc, hepi, hepia, asc40m, ascmc, tchsh1, tchsh2.

Alex was unsure what this additional code was for.  To disable it, we went into the skeleton.st file, and changed the name "SEQUENCER_NAME_lsc" to "SEQUENCER_NAME_lsc_removed" where ever it occured.  These names were in #ifdef statements, so now these codes will only be used if the model is named lsc_removed.  This apparently fixed the problem.  Running startlsc now runs the code as it should, and I can proceed to testing the communication to the lsp model.

Alex said he'd try to figure out what these special #ifdef code pieces are intended for and hopefully completely remove them once we've determined we don't need it.

  2945   Tue May 18 12:04:13 2010 robUpdateIOOFirst steps toward MC mode measuring

Quote:

Another note: Don't trust the PSL shutter and the switch on the MEDM screens! Always use a manual block in addition!!! We discovered upon closeup that hitting the "Closed" button, while it reads back as if the shutter is closed (with the red box around the buttons), does not in fact close the shutter.  The shutter is still wide open.  This must be fixed.

 Has anyone tried pushing the "reset" button on the Uniblitz driver?

  2943   Tue May 18 11:54:29 2010 steveConfigurationSAFETYpsl mechanical shutter is not working

As we learned yesterday, the PSL laser power out put mechanical shutter is not working in the remote mode. It only works in local manual mode.

Do not rely on the MEDM screen monitor readout! The position is only changing on the monitor. The main beam must be blocked before the output periscope.

  2942   Tue May 18 01:40:56 2010 KojiUpdateIOOFirst steps toward MC mode measuring

OK. Don't worry. This is just an initial confusion which we also had for the suspensions a while ago.

The faraday must be clamped. It shakes the table terribly but it is fine. The leveling may change a bit but should be small enough. Otherwise, just tweak the weights. In fact, the faraday has enough large apertures and we hope we don't need to move it again, as far as the MC incident beam is not moved. But if necessary, we don't move the mirrors but move the faraday itself.

Usually the alignment of the MC is taken by MC2/MC3 such that we don't  move the refl. But if you think what have moved is the MC1/MC3 (i.e. activity in the IMC chamber), take the alignment of the MC1/MC3.

It is just a matter of time to get TEM00. If you get TEM11, it is already close. If you align for TEM11, it is enough aligned to lock TEM10 or TEM01. Once you got better mode, align for it again. Eventually you will get TEM00.

The leveling may change by moving the optics and the weight again. But once the leveling is recovered by arranging the weights somewhere else,
the pointing must be fine again. If necessary, You can remove two optics for squeezing injection (strange motorized rotating mirror and a mount sticking out from the table to south.)

Yes, we need to move the PZT mirror. For the connection, only Steve can give us the right way to do it. If it is too much hussle, just move only the mirror and ignore the wiring for now.

I will update how the mirrors should be migrated from the table to the table.

 

  2941   Mon May 17 19:42:11 2010 JenneUpdateIOOFirst steps toward MC mode measuring

[Jenne, Kevin, Steve]

We made some progress toward getting the MC's beam profile measured.  In the end, no changes were made to anything today, but we're more prepared to go for tomorrow.

What we did:

* Grabbed the scanning slit beam scan from the PSL lab.  It's the same kind as we had here at the 40m, so Kevin was able to hook it up to the computer, and confirmed that it works.

* Opened the IOO and OMC chamber doors, and locked the MC.  Unfortunately the MC mode was awful in Yaw.  Awful like TEM(0,10+). But it still locked.  

* Confirmed that the beam went through the Faraday.  I looked at the beam before and after the Faraday on a card, and it was the same nasty beam both before and after.  So it looks like Zach did a good job aligning the Faraday and everything else.  I was going to clamp the Faraday, but I didn't yet, since I wanted to see the nice happy TEM00 mode go through without clipping before risking moving the Faraday during clamping (I don't know how heavy it is, so I'm not sure how much it might potentially move during clamping.)

* Noticed that there is a whole lot of crap on both the OMC and BS tables that's going to have to move.  In particular, one of the weights leveling the OMC table is right where I need to put MMT2.  Steve suggested putting the optic there, in its approximate place, before doing too much other stuff, since it could potentially affect the leveling of the table, and thus the input pointing to the MC.  Unfortunately, to do that I'll need to move the weight, which is definitely going to change things.  Sad face.  Moving the weight will likely be one of the first things I do tomorrow, so that all 3 profile measurements have the same configuration. 

* Before closing up, I tried to align the MC, to get back to TEM00, to no avail.  I got as far as achieving TEM11 flashing, along with a bunch of other crappy modes, but didn't get 00.  That's also on the to-do list.

What we're going to do:

* Open the chambers, and align the MC to TEM00 (using the sliders on the MC align screen).

* Check with an IR card that the beam goes through the Faraday.

* Clamp the Faraday, reconfirm.

* Remove the weight on the OMC table.

* Place MMT2 on the OMC table in it's approximate final location.

* Realign the MC, and make sure the beam goes through the Faraday.  If this doesn't happen smoothly, I may need more instruction since I've never dealt with aligning the Faraday before.  What are the appropriate mirrors to adjust? 

* Move the PZT flat steering mirror from the BS table to the IOO table.  (Thoughts on this?  This will change the table leveling, and also includes the trickiness of needing to move the connectors for the PZT.)

* Place a flat mirror on the BS table to route the MC beam out to the BS/PRM/SRM oplev table. 

* Measure the mode using the beam scan: on the BS oplev table, on the POX table, and then perhaps by shooting the beam through the beamtube on the ETMY (new convention) table.

* Place MMT1 on the BS table, use flat mirrors to get it out of the chambers, repeat measurements.

* Place MMT2 in the correct position, use flat mirrors to get it out of the chambers, repeat measurements.

All of this may require some serious cleaning-up of the BS table, which is going to be ugly, but it has to happen sometime. Hopefully I can get away with only moving a minimal number of things, in order to get these measurements done.

 

Another note: Don't trust the PSL shutter and the switch on the MEDM screens! Always use a manual block in addition!!! We discovered upon closeup that hitting the "Closed" button, while it reads back as if the shutter is closed (with the red box around the buttons), does not in fact close the shutter.  The shutter is still wide open.  This must be fixed.

  2940   Mon May 17 17:17:49 2010 josephb, steve, alberto, kiwamuUpdateCDSNew CDS computers now in racks.

We placed 3 new computers in the racks.  One in 1X4 (machine running SCX) and 2 in 1Y4 (LSC and SUS).  These are 1U chassis, 4 core machines for the CDS upgrade.  I will be bringing over 2 IO chassis and their rails over tomorrow, one to be placed in 1Y4, and 1 in 1X4.

We still need some more 40 pin adapter cables and will send someone over this week to make them.  However, once we have those, we should be able to get two to three machines going, one end computer/chassis and the SUS computer/chassis.

After tomorrow we are still going to be owed 1 computer, another dolphin fiber, a couple of blue boxes, and the LSC, IO, and Y end IO chassis.  We also realized we need further fiber for the timing system.  We're going to need to get and then run fiber to both ends, as well as to 1X3, where the LSC IO chassis will be.

 

  2939   Mon May 17 10:57:16 2010 steveConfigurationSUSdamping restored to ETMYs

ETMY-south sus damping was restored

  2938   Mon May 17 02:10:10 2010 KojiConfigurationIOOHow to lock / align the MC

Let me remind you how to lock and align the IMC

To lock

1. Open the doors for the IMC/OMC chambers. Open the manual shutter of the PSL just in front of the optical window

2. Run scripts/MC/mcloopson

3. Set the MC length path gain 0.3 / Set the MC total gain "+20"

4. If you want to avoid excitation of the mirrors by air turbulence, put a big plastic film and put three posters on the top and both the sides on the floor to block the wind go into the chamber.

To shut down

1. Run scripts/MC/mcloopsoff

2. Close the manual shutter, Remove the wind blockers, and the light door of the chambers

To align the MC

1. Tweak MC2 and MC3 to get maximum transmittion and/or minimum reflection.

  2937   Sun May 16 19:25:45 2010 KojiUpdateGreen Lockingreflected beam at PD

Don't make a short cut. The beam size at a single place does not tell you anything.
Measure the mode of of the beam at multiple points. Calculate the mode matching ratio.

Align the mirrors precisely. Try to see the DC fringe. Predict the size of the DC fringe.

Test the demodulation system with a function generator. Find the 200kHz signal using the spectrum analyzer to find the signal and the optimal alignment.

Put the DC signal and the AC signal to the oscilloscope as X&Y.

Good luck.

 

  2936   Sun May 16 12:51:08 2010 kiwamuUpdateGreen Lockingreflected beam at PD

Mode matching to the cavity has been done.

Now the reflection from the cavity is successfully going into the PD.

However I could not see any obvious error signal.

I should compute and re-check the expected signal level.

 


(mode matching of the crystal)

On the last Wednesday, Kevin and I measured the mode profile before the PPKTP crystal,  and we found the Gaussian beam at the crystal is focused too tightly (w = 38 um).

In order to achieve the best conversion efficiency the waist size should be 50.0 um. So we moved a lens, which was located before the crystal, to 7 cm more away from the crystal. Eventually we obtained a better focus (w = 50.1 um).

Thanks, Kevin. You did a good job.

 

(mode matching of the cavity)

I put a lens with f=-50 mm after the crystal to diverge the green beam more quickly. Then the beam is going through the Faraday of 532 nm, two final modematching lenses and ETMY at last.

By shifting the positions of these lenses, I obtained the reflection from ITMY with almost the same spot size as that of the incident. This means modemathing is good enough.

I put two more steering mirrors before its injection to the ETM, this allows us to align the beam axis against the cavity.

I aligned the axis by using the steering mirrors and now the green beam are successfully hitting the center of both the ETM and the ITM.

Then the alignment of the ETM and the ITM was adjusted from medm, so that both reflection goes in the same path as that of the incident.

And then I put a PD (Thorlabs PDA36A) to see the reflection rejected by the Faraday.

Connecting a mixer and a local oscillator (Stanford func. generator) with f=200kHz, but I couldn't see any obvious PDH signal....

Since the PD is DC coupled, the signal is almost dominated by DC voltage. Even if I inserted a high pass filter to cut off the DC, the AC signal looks very tiny..

  2935   Sat May 15 04:13:33 2010 KojiSummaryIOOMC table leveled

Fixing at the next time is absolutely OK.

Quote:

Ah... no, I didn't. That explains why there were loose dogclamps on the table. I wrapped them in foil and put them on the clean cart. Can this wait until the next time we open the tank (i.e. to measure the beam profile), or should I go over there and clamp it down today? 

 

  2934   Fri May 14 16:19:22 2010 JenneUpdatePEMGuts of a Guralp

[Jenne, Rana]

We took apart and examined one of the Guralp seismometers this afternoon.  For the most part we think we understand how it works. The horizontal sensors are a little more confusing, since we didn't end up finding the moving masses.  The vertical sensor is a flat rectangle, hinged at one edge.  There are capacitive sensors above and below the rectangle.  The hinged end is connected to a leaf spring. 

The PCBs are packed full of old-school 80's components.  We probably need an actual schematic to figure out where the preamp circuit is, which is what we'd want to think about fitzing with, if we were to try to improve the noise of the seismometer.  For now, we put it all back together, and back out on the granite slab. 

There was a wee bit of confusion when putting the N/S marker-spikes back on as to where they should go.  The solution is that the handle of the seismometer is aligned with the North/South axis, so the spikes should be aligned with the handle.  The lid of the seismometer is uniquely aligned to the stuff inside by the ribbon cable connector, as well as the holes in the lid for accessing the centering potentiometers.  So, align the lid to the pots, and then align the spikes to the handle.

Photos are on Picasa.

  2933   Fri May 14 16:14:37 2010 KevinUpdateGreen LockingGreen Laser Beam Profile

Quote:

Strange. I thought the new result became twice of the first result. i.e. w0=32um or so.

Can you explain why the waist raidus is estimated to be three times of the last one?
Can you explain why the measured radius @~70mm is not 0.8mm, which you told us last time,
but is 0.6mm?

The measurements have been done at the outside of the Rayleigh range.
This means that the waist size is derived from the divergence angle

theta = lambda / (pi w0)

At the beginning you used diameter instead of radius. This means you used twice larger theta to determine w0.
So if that mistake is corrected, the result for w0 should be just twice of the previous wrong fit.

 

 

I was off by a factor of sqrt(2). The correct fit parameters are

for the vertical beam profile:

reduced chi^2 = 3.28

x0 = (-87 ± 1) mm

w0 = (32.59 ± 27) µm

for the horizontal beam profile

reduced chi^2 = 2.02

x0 = (-82 ± 1) mm

w0 = (32.23 ± 20) µm

In the following plots * denotes vertical data points and + denotes horizontal data points. The blue curve is the fit to the vertical data and the purple curve is the fit to the horizontal data.

Attachment 1: profile.png
profile.png
Attachment 2: residuals.png
residuals.png
  2932   Fri May 14 12:14:26 2010 josephbUpdateCDSNeed to track down old code for lsc system and remove them

I'm currently in the process of tracking down what legacy code is interfering with the new lsc model.

It turns out if you change the name of lsc file to something else (say scx as a quick test for example), it runs fine.  In fact, the lsc and scx GDS_TP screens work in that case (since they're looking at the same channels).  As one would expect, running them both at the same time causes problems.  Note to self, make sure the other one is killed first.  It does mean the lsc code gets loaded part way, but doesn't seem to communicate on EPICs or to the other models.  However, I don't know what existing code is interfering.  Currently going trhough the target directories and so forth.

  2931   Fri May 14 10:33:01 2010 ZachSummaryIOOMC table leveled

Ah... no, I didn't. That explains why there were loose dogclamps on the table. I wrapped them in foil and put them on the clean cart. Can this wait until the next time we open the tank (i.e. to measure the beam profile), or should I go over there and clamp it down today?

 

  2930   Fri May 14 08:18:46 2010 steveUpdateGreen LockingReflection from ETM and ITM !

I stopped AWG  1 Hz drive to ITMYs. ITMXe was also driven or oscillating. ITMXe damping was off, so I turned it on. It did not effect it's oscillation

Attachment 1: itmx1hzos.jpg
itmx1hzos.jpg
  2929   Fri May 14 03:30:45 2010 KojiSummaryIOOMC table leveled

Thanks Zach.This was a great job.

It was not mentioned but: was the Faraday clamped down on the table?

 

  2928   Thu May 13 23:59:46 2010 ZachSummaryIOOMC table leveled

 After the recent removal of the old IMMT and the relocation of the Faraday isolator, the MC table was tilted a bit (southward and slightly westward---as of when I opened the chamber this afternoon). I re-leveled it by putting an extra two rectangular ballast blocks on the stack that was already hanging off the NNE edge of the table (there are a total of 4 in the stack now). I also screwed down the circular block that Koji and I put between the Faraday and SM1 on Tuesday, and re-mounted the two wire harness towers onto the table.

Needless to say, this threw the MC way out of alignment. I spent the rest of the afternoon reacquiring alignment and getting it to lock robustly. Here is a summary:

  • I adjusted MC3 until I got the 2nd, 3rd+ pass beams to overlap with the input beam between MC1&3, then I adjusted MC2&1 semi-methodically until I got something flashing at the transmitted end. This took some time.
  • I went back into the control room, engaged the loops and acquired lock on the TEM00 mode, whereupon I found that the beam spot was WAY off center on MC2 (due to my meddling with all the mirrors to get resonance flashes). I began using the MC2_spot_up (etc) scripts we wrote the other day to re-center it.
  • After a few iterations, the lock became weak, and eventually gave out. This is because the REFL beam was falling off the RFPD (and being clipped by the iris on the AP table), so I moved the iris and re-centered the beam on the diode.
  • With that, I was able to get the MC2 spot more or less centered, but then I noticed that---though the lock was clearly strong as evidenced both by the REFL power dip and visually via the camera on MC2---it looked like crap on the CCD. It seemed like there was some higher order mode structure sloshing around on top of the 00 spot, which didn't make any sense, until I realized that it was just a diffraction pattern from the TRANS beam getting clipped somewhere on the way out of the vacuum system.
  • I went back to the AP table, where I noticed that the TRANS beam was hitting near the edges of several of the mirrors on the way back to the PSL table, including the first one out of the viewport, so I turned IM4 to center the beam on this mirror, then proceeded to center the beam on each mirror downstream and then onto the CCD.
  • After getting a clear picture of the transmission on the CCD, centering the spot even better on MC2, then fine-tuning MC2&3 to strengthen the lock, I went back to the MC table to check that the transmitted beam was still passing through the center of the Faraday, which, by none other than an act of God, it was.
  • Having done the necessary work in the tank, I ran the A2L_MC2 script to fine-tune the centering of the spot on MC2. It needed a couple steps up and to the side, but after that the actuator gains for pitch and yaw were both balanced again to within ~2%, which is only slightly above the measurement error. We will probably need to adjust this continually, especially during the upgrade, so I didn't bother with getting it better than that.

After that, I shut off the loops, blocked the beam, and put the light doors back on the tanks. Then I went to the parking lot, then I got in my car, etc, etc, etc.

  2927   Thu May 13 15:19:44 2010 josephbUpdateCDSTrying to get lsc.mdl and lsp.mdl working

I had a chat with Alex this morning and discovered that the dcu_ids 13,14,15,16 are reserved currently, and should not be used.  I was told 9-12 and 17-26 were fine to use.  I pointed out that we will eventually have more modules than that.  His response was he is currently working on the framebuilder code and "modernizing" it, and that those restrictions will hopefully be lifted in the future although he isn't certain at this time what the real maximum gds_id number is (he was only willing to vouch for up to 26 - although the OMC seems to be currently working and set to 30).

Alex also suggested running an iop module to provide timing (since we are using adcSlave=1 option in the models).  Apparently these are x00.mdl, x01.mdl, x11.mdl files in the /home/control/cds/advLigoRTS/src/epics/simLink/ directory.  I saved x00.mdl as io1.mdl (I didn't want to use io0 as its a pain to differentiate between a zero and 'O'.  This new IOP is using gds_node=1, dcu_id=9.  I modified the approriate files to include it.

I modified /etc/rc.d/rc.local and added io1 to shmem line.  I modified /cvs/cds/caltech/target/fb/daqdrc to use dcu_id 9 as the controller (this is the new iop model dcu_id number).  In that same directory I modifed the file master by adding /cvs/cds/caltech/chans/daq/C1IO1.ini as well as uncommenting tpchn_C1 line.  I modified testpoint.par in /cvs/cds/caltech/target/gds/param to include C-node0, and modified the prognum for lsc and lsp to 0x31001003 and 0x31001005.

So I started the 3 processes with startio1, startlsc, startlsp, then went to the fb directory and started the framebuilder.  However, the model lsc.mdl is still having issues, although lsp and io1 seem to be working.  At this point I just need to track down what fundamentally is different between lsc and lsp and correct it in the lsc model.  I'm hoping its not related to the fact that we actually had a previous lsc front end and there's some legacy stuff getting in the way.  One thing I can test is changing the name and see if that runs.

 

  2926   Thu May 13 05:06:43 2010 ranaUpdate40m Upgrading216 MHz resonance in the POY11 PD killed

 

 This idea was tried before by Dale in the ~1998 generation of PDs. Its OK for damping a resonance, but it has the unfortunate consequence of hurting the dynamic range of the opamp. The 100 Ohm resistor reduces the signal that can be put out to the output without saturating the 4107.

I still recommend that you move the notch away from the input of the 4107. Look at how the double notch solution has been implemented in the WFS heads.

  2925   Wed May 12 23:31:17 2010 AlbertoUpdate40m Upgrading216 MHz resonance in the POY11 PD killed
It turned out that the resonance at 216 MHz in the 11MHz PD that I showed in the elog entry 2902 was casued by an instability of the of the MAX4107 opamap' feedback loop.
As the datasheet of the opamp shows, the close-loop gain has a peak at about 200-250MHz, in presence of even small capacitive loads.
In my case, perhaps either the capacitance of the BNC cables plugged to the RF output of the PD box, or the shunt capacitance of the circuit parts after the opamap (traces and resistors) might have introduced capacitance at the output of the amplifier.
 
LISO had failed in predicting the resonance because it has only ideal transfer functions of the opamps. In particular the open-loop gain of the opamaps in the library is just a function with a simple pole.
 
At RF frequencies the output impedances of the opamp starts having a non-negligible inductance that interacts with the load capacitance, generating a typical LC-circuit resonance.
 
In cases like this, such effect can be mitigated by introducing an "isolating" resistor at the output of the opamp.
 
So I did that and modified the circuit as in this simplified schematic here:
 
POX11.png
 

The choice of 100 Ohm for the isolating resistor was mainly empirical. I started with 10, then 20 and 50 until I got a sufficient suppression of the resonance. Even just 10Ohm suppressed the resonance by several tens of dB.

2010-05-12_POY11_CalibratedOpticalResponse0-500MHz.png

 

In that way the gain of the loop didn't change. Before that, I was also able to kill the resonance by just increasing the loop gain from 10 to 17.  But, I didn't want to increase the closed-loop gain.

One thing that I tried, on Koji's suggestion, was to try to connect the RF output of the PD box to an RF amplifier to see whether shielding the output from the cable capacitance would make the resonance disappear: It did not work.

  2924   Wed May 12 17:10:16 2010 AlbertoUpdate40m UpgradingRF frequency generation box - step 0

I started putting together the components that are coint to go inside the frequency generation box. Here's how it looked like:

DSC_1499_small.JPG

The single component are going to be mounted on a board that is going to sit on the bottom of the box.

I'm thinking whether to mount the components on an isolating board (like they did in GEO), or on an aluminum board.

I emailed Hartmut to know more details about his motivations on making that choice.

  2923   Wed May 12 12:58:26 2010 josephbConfigurationCDSSetup fb to handle lsc, lsp models on megatron

I modified /cvs/cds/caltech/target/fb and changed the line "set controller_dcu=10" to "set controller_dcu=13" (where 13 is the lsc dcu_id number).

I also changed the set gds_server line from having 10 and 11 to 13 and 14 (lsc and lsp).

The file /cvs/cds/caltech/fb/master was modified to use C1LSC.ini and C1LSP.ini, as well as tpchn_C2.par (LSC) and tpchn_C3.par (LSP)

testpoint.par in /cvs/cds/caltech/target/gds/param was modified to use C-node1 and C-node2 (1 less then the gds_node_id for lsc and lsp respectively).

Note all the values of gds_node_id, dcu_id, and so forth are recorded at http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/Existing_RCG_DCUID_and_gds_ids

  2922   Wed May 12 12:32:04 2010 josephbConfigurationCDSModified /etc/rc.d/rc.local on megatron

I modified the /etc/rc.d/rc.local file on megatron removing a bunch of the old test module names and added the new lsc and lsp modules, as well as a couple planned suspension models and plants, to shared memory so that they'll work.  Basically I'm trying to move forward into the era of working on the actual model we're going to use in the long term as opposed to continually tweaking "test" models.

The last line in the file is now: /usr/bin/setup_shmem.rtl lsc lsp spy scy spx scx sus sup&

I removed mdp mdc mon mem grc grp aaa tst tmt.

  2921   Wed May 12 12:25:11 2010 KojiUpdateGreen LockingRe: Reflection from ETM and ITM !

??? I still don't understand. What principle are you rely on?

I could not understand why you rotated the HWP to the "minimum" transmission
and then minimized the transmission by rotating the output PBS. What is optimized by this action?

Probably there is some hidden assumption  which I still don't understand.
Something like:
Better transmission gives best isolation, PBS has some leakage transmission
of the S-pol light, and so on.

Tell me what is the principle otherwise I don't accept that this adjustment is "to get a good isolation with the Faraday".

P.S. you could flip the faraday without removing it from the V-shaped mount. This does not roll the Faraday.

Quote:

The procedure you wrote down as a standard is right.   I explain reasons why we didn't do such way. 

For our situation, we can rotate the polarization angle of the incident beam by using a HWP in front of the Faraday.  

This means we don't have to pay attention about the PBS_in because the rotation of either PBS_in or the HWP causes the same effect (i.e. variable transmission ). This is why we didn't carefully check the PBS_in, but did carefully with the HWP.

Normally we should take a maximum transmission according to a instruction paper from OFR, but we figured out it was difficult to find a maximum point. In fact looking at the change of the power with such big incident (~1W) was too hard to track, it only can change 4th significant digit ( corresponds to 1mW accuracy for high power incident ) in the monitor of the Ophir power meter. So we decided to go to a minimum point instead a maximum point, and around a minmum point we could resolve the power with accuracy of less than 1mW.

After obtaining the minimum by rotating the HWP, we adjusted the angle of PBS_out to have a minimum transmission.

And then we was going to flip the Faraday 180 deg for fine tuning, but we didn't. We found that once we remove the Faraday from the mount, the role angle of the Faraday is going to be screwed up because the mount can not control the role angle of the Faraday. This is why we didn't flip it.

Quote:

I could not understand this operation. Can you explain this a bit more?

It sounds different from the standard procedure to adjust the Faraday:

1) Get Max transmittion by rotating PBS_in and PBS_out.

2) Flip the Faraday 180 deg i.e. put the beam from the output port.

3) Rotate PBS_in to have the best isolation.

 

 

 

  2920   Wed May 12 10:33:32 2010 kiwamuUpdateGreen LockingRe: Reflection from ETM and ITM !

The procedure you wrote down as a standard is right.   I explain reasons why we didn't do such way. 

For our situation, we can rotate the polarization angle of the incident beam by using a HWP in front of the Faraday.  

This means we don't have to pay attention about the PBS_in because the rotation of either PBS_in or the HWP causes the same effect (i.e. variable transmission ). This is why we didn't carefully check the PBS_in, but did carefully with the HWP.

Normally we should take a maximum transmission according to a instruction paper from OFR, but we figured out it was difficult to find a maximum point. In fact looking at the change of the power with such big incident (~1W) was too hard to track, it only can change 4th significant digit ( corresponds to 1mW accuracy for high power incident ) in the monitor of the Ophir power meter. So we decided to go to a minimum point instead a maximum point, and around a minmum point we could resolve the power with accuracy of less than 1mW.

After obtaining the minimum by rotating the HWP, we adjusted the angle of PBS_out to have a minimum transmission.

And then we was going to flip the Faraday 180 deg for fine tuning, but we didn't. We found that once we remove the Faraday from the mount, the role angle of the Faraday is going to be screwed up because the mount can not control the role angle of the Faraday. This is why we didn't flip it.

Quote:

I could not understand this operation. Can you explain this a bit more?

It sounds different from the standard procedure to adjust the Faraday:

1) Get Max transmittion by rotating PBS_in and PBS_out.

2) Flip the Faraday 180 deg i.e. put the beam from the output port.

3) Rotate PBS_in to have the best isolation.

 

 

  2919   Wed May 12 09:16:29 2010 steveUpdateGreen LockingReflection from ETM and ITM !

 

 Now I know why Rana was wearing his bright green pants yesterday. It is nice to see the green beam in the 40m IFO again. It calls for celebration!

I stopped AWG 1Hz drive of ITMYs (south-arm) I still see unblocked beams at the ETMYs table. We have plenty of cleaned razor beam traps to be used. Please block Faraday rejects etc

  2918   Wed May 12 03:56:54 2010 KojiUpdateIOOFaraday aligned

Zach and Koji

The old small MMT was removed and wrapped by Al foils.

The steering mirror IM2-IM4 were displaced and aligned.

The Faraday isolator block is moved and aligned.

The MC is realigned and resonatng TEM-00.

Now the MC has slightly miscentered beam on the mirrors owing to change of the stack leveling.
OSEMs are also in a strange state. We should check this later.

  2917   Wed May 12 03:52:54 2010 KojiUpdateGreen LockingReflection from ETM and ITM !

I could not understand this operation. Can you explain this a bit more?

It sounds different from the standard procedure to adjust the Faraday:

1) Get Max transmittion by rotating PBS_in and PBS_out.

2) Flip the Faraday 180 deg i.e. put the beam from the output port.

3) Rotate PBS_in to have the best isolation.

Quote:

* To get a good isolation with the Faraday we at first rotated the polarization of the incident beam so to have a minimum transmission. And then we rotated the output polarizer until the transmission reaches a minimum. Eventually we got the transmission of less than 1mW, so now the Faraday should be working regardless of the polarization angle of the incident beam. As we predicted, the output polaerizer seems to be rotated 45 deg from that of the input.

  2916   Wed May 12 03:42:38 2010 KojiUpdateGreen LockingGreen Laser Beam Profile

Strange. I thought the new result became twice of the first result. i.e. w0=32um or so.

Can you explain why the waist raidus is estimated to be three times of the last one?
Can you explain why the measured radius @~70mm is not 0.8mm, which you told us last time,
but is 0.6mm?

The measurements have been done at the outside of the Rayleigh range.
This means that the waist size is derived from the divergence angle

theta = lambda / (pi w0)

At the beginning you used diameter instead of radius. This means you used twice larger theta to determine w0.
So if that mistake is corrected, the result for w0 should be just twice of the previous wrong fit.

Quote:

 

I recalculated the fits using the radius of the beam instead of the diameter of the beam at 13.5% full-width with the following results:

For the vertical beam profile:

reduced chi^2 = 3.25

x0 = (-86 ± 1)mm

w0 = (46.01 ± 0.38)µm

For the horizontal beam profile:

reduced chi^2 = 2.05

x0 = (-81 ± 1)mm

w0 = (45.50 ± 0.28)µm

 

  2915   Wed May 12 02:35:13 2010 Koji, Rana, KiwamuUpdateGreen LockingReflection from ETM and ITM !

We succeeded in getting the reflected green beam from both ITMY and ETMY.

After we did several things on the end table, we eventually could observe these reflections.

Now the spot size of the reflection from ITMY is still big ( more than 1 cm ), so tomorrow modematching to the 40m cavity is going to be improved by putting mode matching telescopes on right positions.

An important thing we found is that, the beam height of optics which directly guides the beam to the cavity should be 4.5 inch on the end table.

 


(what we did)

* Aidan, Kevin and Kiwamu set the beam to be linearly polarized by rotating a QWP in front of the Innolight. This was done by monitoring the power of the transmitted light from the polarizer attached on the input of the Faraday of 1064 nm. Note that the angle for QWP is 326.4 deg.

* We put some beam damps against the rejected beam from the Faraday

* To get a good isolation with the Faraday we at first rotated the polarization of the incident beam so to have a minimum transmission. And then we rotated the output polarizer until the transmission reaches a minimum. Eventually we got the transmission of less than 1mW, so now the Faraday should be working regardless of the polarization angle of the incident beam. As we predicted, the output polaerizer seems to be rotated 45 deg from that of the input.

* Rana, Koji and Kiwamu aligned the PPKTP crystal to maximize the power of 532 nm.  Now the incident power of 1064 nm is adjusted to 250mW and the output power for 532 nm is 0.77mW. Actually we can increase the laser power by rotating a HWP in front of the Faraday.

* We injected the green beam to the chamber and aligned the beam axis to the ETMY without the modematching lenses, while exciting the horizontal motion of the ETM with f=1Hz from awg. This excitation was very helpful because we could figure out which spot was the reflection from the ETM.

* Once we made the reflected beam going close to the path of the incident beam, we then put the modematching lenses and aligned the steering mirrors and lenses. At this time we could see the reflected beam was successfully kicked away by the Faraday of 532 nm.

* Koji went to ITMY chamber with a walkie-talkie and looked at the spot position. Then he told Rana and Kiwamu to go a right direction with the steering mirrors. At last we could see a green beam from ITM illuminating the ETM cage.

* We excited the ITMY with f=2Hz vertically and aligned the ITM from medm. Also we recovered a video monitor which was abandoned around ETMY chamber so that we could see the spot on the ETM via the monitor. Seeing that monitor we aligned the ITM and we obtained the reclection from the ITM at the end table.

* We also tried to match the mode by moving a lens with f=400mm, but we couldn't obtain a good spot size.

 

ELOG V3.1.3-